OpenStudy (anonymous):

can someone pls solve this equation: F=9/5 c+32 F=9/5 (325) +32 and im lost from here any help with this problem i thnk u

OpenStudy (zzr0ck3r):

c = 325 notice that is the only thing that will make these the same

OpenStudy (zzr0ck3r):

wait im confused. (9/5)(325) + 32 = 9*65+32 ....is that what your asking?

OpenStudy (anonymous):

um i think...im trying to figured out how u set the problem up to be solved using the system F=9/5 C+32 Fis for fahrenheit and C is degrees Celsius

OpenStudy (zzr0ck3r):

and they gave you f or c?